# MiLe Loss: a New Loss for Mitigating the Bias of Learning Difficulties in Generative Language Models

### 21 Feb 2024 v5

@@ 1 Introduction
It is a straightforward idea to use the notable Focal Loss Lin et al. (2017) from the field of object detection as an alternative to the prevalent Cross-Entropy Loss for the next token prediction. This modification aims to intensify the language model’s focus on the infrequent and difficult-to-learn tokens. Focal Loss is a dynamically scaled version of the Cross-Entropy Loss, where the scaling factor decreases as the predicted probability w.r.t the ground-truth token increases. Specifically, Focal Loss decreases the weights of the easy-to-learn tokens, as their predicted probabilities are higher, and meanwhile increases the weights of the difficult-to-learn ones, as their predicted probabilities are lower. In that way, it compels the language model to pay more attention to difficult-to-learn tokens.
Nevertheless, Focal Loss Lin et al. (2017) only takes into account the probability w.r.t the ground-truth token when assessing its learning difficulty, and is intuitively designed for the multi-class classification problem where an object is only associated with a single class label. Indeed, in language model pretraining, when predicting the next token given the previous ones, there might exist multiple valid tokens besides the ground-truth one. This makes predicting the next token more like a multi-label classification problem, where an object can be associated with multiple class labels. For example, as shown in Figure 1, given the previous tokens “I like playing ”, there are multiple valid next tokens, like “basketball”, “football”, “golf”, etc. Suppose the target training token sequence is “I like playing basketball”. As the valid tokens would divide up almost the total probability (i.e., 1.0), the ground-truth token “basketball” would be given a smaller probability (e.g., 0.18). Then Focal Loss would treat “basketball” for the position as a difficult-to-learn token. However, as all the other valid tokens are also correct for the position in the view of language modeling, only allowing “basketball” to be predicted is unsuitable. Thus, the learning difficulty assessed by Focal Loss for “basketball” is imperfect in such a multi-label classification case.

### 25 Mar 2024 v6

@@ 1 Introduction
We propose to utilize the notable Focal Loss Lin et al. (2017) from the field of object detection as an alternative to the prevalent Cross-Entropy Loss for the next token prediction. This modification aims to intensify the language model’s focus on the infrequent and difficult-to-learn tokens. Focal Loss is a dynamically scaled version of the Cross-Entropy Loss, where the scaling factor decreases as the predicted probability w.r.t the ground-truth token increases. Specifically, Focal Loss decreases the weights of the easy-to-learn tokens, as their predicted probabilities are higher, and meanwhile increases the weights of the difficult-to-learn ones, as their predicted probabilities are lower. In that way, it compels the language model to pay more attention to difficult-to-learn tokens.
Nevertheless, Focal Loss Lin et al. (2017) only takes into account the probability w.r.t the ground-truth token when assessing its learning difficulty, and is intuitively designed for the multi-class classification problem where an object is only associated with a single class label. Indeed, in language model pretraining, when predicting the next token given the previous ones, there might exist multiple valid tokens besides the ground-truth one. This makes predicting the next token more like a multi-label classification problem, where an object can be associated with multiple class labelsTsoumakas and Katakis (2007); Chen et al. (2018). For example, as shown in Figure 1, given the previous tokens “I like playing ”, there are multiple valid next tokens, like “basketball”, “football”, “golf”, etc. Suppose the target training token sequence is “I like playing basketball”. As the valid tokens would divide up almost the total probability (i.e., 1.0), the ground-truth token “basketball” would be given a smaller probability (e.g., 0.18). Then Focal Loss would treat “basketball” for the position as a difficult-to-learn token. However, as all the other valid tokens are also correct for the position in the view of language modeling, only allowing “basketball” to be predicted is unsuitable. Thus, the learning difficulty assessed by Focal Loss for “basketball” is imperfect in such a multi-label classification case.
@@ 5.2 Perplexity on the Pile Validation Set
BoolQ
HellaSwag
LAMBADA
OpenBookQA
PIQA
SIQA
StoryCloze
Winogrande
Avg
0-shot Cross-Entropy Loss
66.73
63.48
60.95
36.80
75.52
44.58
72.26
61.88
60.27
MiLe Loss 68.62 64.17 61.52 39.00
75.41
44.63 72.90 63.61 61.23
1-shot Cross-Entropy Loss
64.13
63.33
58.92
40.40
75.46
48.72
72.90
63.85
60.96
MiLe Loss 65.26 63.93 60.57
38.80
75.46 49.64
72.58
63.93 61.27
5-shot Cross-Entropy Loss
64.22
63.92
60.90
39.60
75.84
51.18
73.60
64.72
61.75
MiLe Loss 66.85 64.58 64.33 41.00
75.14
52.66 74.02 66.06 63.08
Table 8: The performance of the 6.7B models trained with 200B tokens in zero/few-shot settings across various benchmarks.
5.3 Performance of Training with More Tokens
MiLe Loss assesses the learning difficulty of each token through information entropy. Intuitively, the more tokens used in model training, the more powerful the Language Model (LM) becomes, and the output word distribution becomes more reasonable. Consequently, the assessment of the learning difficulties of tokens becomes more accurate, and thus the MiLe Loss can probably better lead the LM to tackle the bias. To validate that, with limited computational resources, we continue to pre-train the 6.7B model from 100B tokens to 200B tokens, with both Cross-Entropy Loss and MiLe Loss. Their corresponding evaluation results on all benchmarks are reported in the following Table 8. We can see that, when the number of training tokens for the 6.7B models increases to 200B, the models trained with MiLe Loss yield consistent and substantial performance improvements over those trained with Cross-Entropy Loss. Moreover, compared to training with 100B tokens, training with more tokens even helps MiLe Loss to yield LARGER performance improvements. For instance, in the 5-shot setting, with 100B training tokens, the performance improvements gained by MiLe Loss over Cross-Entropy Loss on the 6.7B models is 1.02%. Then by continuing pre-training with more training tokens, the gained improvements increase to 1.33%. The experimental results above demonstrate well that using more tokens and longer pre-training increases the benefits of MiLe Loss.